OpenStudy (anonymous):

5/x-1 + 3/x+1 = -4x/x^2 - 1

OpenStudy (saifoo.khan):

\[\frac{5}{x-1}+\frac{3}{x+1}= -\frac{-4x}{x^2-1}\]this?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

\[x = -\frac{1}{6}\]

OpenStudy (anonymous):

LCM of (x-1)and (x+1) =(x-1)*(x+1) now u can easily solve this \[5/(x-1) + 3/(x+1)\] let us forget the right hand side ok now \[[5(x+1)+3(x-1)]/(x-1)(x+1)\] \[(5x+5+3x-3 )/(x^2-1)\] \[(8x+2)/(x^2-1)\] now it is the turn of right hand side transpose it to left hand side \[(8x+2)/(x^2-1)\] +\[4x/(x^2-1)\] that is equal to \[(8x+2+4x)/(x^2-1)\] =0 \[(12x+2)/(x^2-1)\] = 0 12x+2= 0 12x=-2 x=12/(-2) x=-6

OpenStudy (saifoo.khan):

kush, u r wrong.

OpenStudy (anonymous):

yes sorry i did wrong between

OpenStudy (anonymous):

the answer is -1/2 (8x+2-4x)/(x^2-1)=0 4x+2=0 4x=-2 x=-1/2

OpenStudy (anonymous):

no it says thats the wrong answer

OpenStudy (saifoo.khan):

Mine is right! otherwise check the equation.

OpenStudy (anonymous):

urs is right saif

OpenStudy (saifoo.khan):

I knew! :D

OpenStudy (anonymous):

:D